### v4

@@ Abstract
Large language models have gained significant traction and popularity in recent times, extending their usage to code-generation tasks. While this field has garnered considerable attention, the exploration of testing and evaluating the robustness of code generation models remains an ongoing endeavor. Previous studies have primarily focused on code generation models specifically for the Python language, overlooking other widely-used programming languages. In this work, we conduct a comprehensive comparative analysis to assess the robustness performance of several prominent code generation models and investigate whether robustness can be improved by repairing perturbed docstrings using an LLM. Furthermore, we investigate how their performance varies across different programming languages. To accomplish this, we introduce perturbations in four key areas of the prompt: DocString, function name, syntax, and format. We have compiled and released a dedicated dataset for this purpose. This work presents our experimental findings, shedding light on the performance of code generation models in various scenarios.
@@ 1 Introduction
In this study, we conduct a cross-language robustness evaluation of large language models for code generation across Java, C++, and JavaScript, aiming to provide a broader understanding of how LLMs behave when exposed to perturbed prompts in different programming languages. We evaluate six LLMs on nominal and perturbed datasets, measuring how robustness varies under different perturbation types and identifying language-dependent differences in performance.
We also introduce a mitigation approach that repairs perturbed docstrings using a secondary LLM prompt prior to code generation. This repair strategy is applied using Magicoder, chosen for its ability to interpret natural-language documentation. The results show slight robustness improvements for several perturbation types, with occasional degradation in others, revealing both the promise and the limitations of lightweight prompt-level repair techniques.
@@ 2.1 Background
In this work, we focus on evaluating the robustness of code generation models. During robustness testing, we apply small, semantics-preserving modifications to the original prompt, for example, changing a single character in a function name, and then ask the model to generate code from this perturbed input. After generation, the correctness of the model’s output is assessed using the same test cases as in the nominal (unperturbed) setting. Following prior work, ReCode Wang et al. (2022), which studies robustness only in Python, we adapt and extend its 29 perturbation techniques to support C++, Java, and JavaScript.
@@ 2.2 Related Work
Besides, recent studies have explored various aspects of code generation. Such as, Li et al. (2024a) and Li et al. (2024b) have focused on secure code generation, while Ling et al. (2025) investigate social bias in code generation. Lin et al. (2024) examine code generation through the lens of emulating the software development process, and Saha et al. (2024) propose methods for generating code using the source code from other programming languages.
@@ 3.2 Dataset Preparation
To address the limitations of Wang et al. (2022), which only evaluates code generation robustness in a single programming language (i.e., Python), we extend the evaluation to a multilingual setting by utilizing the HumanEval-X dataset. HumanEval-X, introduced by Zheng et al. (2023), serves as a benchmark for assessing the multilingual capabilities of code generation models. The dataset includes Python, C++, Java, JavaScript, and Go. We exclude Go because it is less commonly used compared to the other languages in both research and industrial contexts. We also exclude Python from our evaluation since it was already thoroughly studied in prior work, ReCode Wang et al. (2022). We focused on C++, Java, and JavaScript, which represent diverse language paradigms (compiled vs. interpreted, statically vs. dynamically typed) and are widely used in real-world software development. This selection allows for a balanced and representative evaluation across different programming styles.
@@
We extend this dataset to multilingual by manually rewriting the prompts, canonical solutions, and test cases in C++, Java, and JavaScript through the following steps:
@@
    We find it necessary to update certain canonical solutions because they failed to handle corner cases, inputs at the edges of valid ranges or uncommon conditions that reveal latent bugs. Figures 2 and 8 show two representative failures: in Figure 2 the model produces a wrong answer for a corner input, while in Figure 8 the original solution leads to a timeout. Both examples demonstrate that passing the standard HumanEval-X tests does not guarantee robustness. To mitigate this, we follow the EvalPlus methodology to refine the affected canonical solutions manually by ourselves so they correctly handle these edge cases. After applying those fixes, we generate prompts from the revised solutions for subsequent evaluation.
@@
    We convert the test cases from EvalPlus, originally written in Python, into Java, C++, and JavaScript. During this process, we ensure that data types are transformed into their equivalent counterparts (e.g., converting lists to ArrayList in Java and lists to vector in C++).
@@
    We exclude test cases that are not compatible with Java and C++ (e.g., cases with mixed lists of integers, strings, and lists). Additionally, we attempt to modify the existing canonical solutions in Java, C++, and JavaScript to accommodate the converted test cases. However, if the updated solutions are unable to handle the test cases effectively, we exclude those specific test cases from the evaluation. For instance, Python allows for unlimited integer sizes, and some EvalPlus test cases tailored for Python might have very large integer values, surpassing the capacity of data types used in Java, C++, and JavaScript. Despite this challenge, we strive to maintain the method signature, including return types and parameter types, identical to those found in the original prompts.
@@ 4.1 Dataset
EvalPlus. EvalPlus Liu et al. (2023) is an enhanced version of HumanEval, originally developed for Python. It introduces a substantially larger set of corner cases and refined prompts that are better aligned with the associated test cases. In this work, we adapt the EvalPlus test cases for our three target programming languages, following the curation process described in Section 3.2. The prompts are largely consistent with those in HumanEval-X, except in instances where minor modifications were required to accommodate language-specific corner cases.
@@ 4.3 Setup
The same set of hyperparameters is applied across models, chosen through a small grid search on a held-out subset of the evaluation data. The temperature is set to 0.2 to reduce randomness and increase the likelihood of generating syntactically valid and test-passing code. The nucleus sampling parameter (top_p) is set to 0.95 to retain diversity while suppressing low-probability outputs. These values are taken from the ReCode Wang et al. (2022). The maximum generation length is set to 1536 tokens to prevent truncation. This value is based on our experience that it provides sufficient output while keeping inference time low.
@@ 5.1 RQ1: How robust are LLMs across different programming languages?
Approach: To investigate this research question, we conduct experiments on Java, C++, and JavaScript using 2 datasets: 1st is HumanEval-X, and 2nd is our curated dataset derived from HumanEval-X Zheng et al. (2023) and EvalPlus Liu et al. (2023). We evaluate six LLMs for code generation: Incoder-1B, Incoder-6B, CodeGen-Multi-2B, CodeGen-Multi-6B, Magicoder-7B, and QwenCode-3.5-6.7B. We evaluate the performance using three metrics: $RP_{s}@k$ measures robustness, $RD_{s}@k$ assesses the decrease in robustness due to changes, and $RR_{s}@k$ captures the relative differences between the original and modified datasets. The perturbations related to the Function Name and the DocString are compared with the nominal prompts (prompts that have only DocString and method signature), while the perturbations related to Syntax and Format are compared with the partial prompts (prompts that have some lines of code along with DocString and method signature).
@@
    Stronger test cases expose more failures across all languages. From Tables 8 and 9, it is evident that the stronger test cases from EvalPlus reveal substantially more failures in both nominal and perturbed settings, which aligns with prior findings Liu et al. (2023). The same trend holds for the partial cases, as shown in Tables 10 and 11. However, the robustness drop remains roughly consistent regardless of the absolute performance levels.
@@ 5.3 Can the robustness drop be mitigated by fixing the perturbed prompts?
Approach: We focus on the DocString perturbation dataset for this experiment. Our approach is inspired by prior work in natural language processing, where LLMs are used to correct typographical and grammatical errors in English sentences Coyne et al. (2023). Specifically, we select all ten DocString-perturbed subsets from our dataset and employ the Magicoder-7B model to fix the prompts. We choose Magicoder because it is instruction-tuned and capable of performing multi-purpose tasks among our evaluated models.
@@
Prompt for Fixing Perturbed DocStrings fix the grammatical errors, spelling, and typos in the following text and reply with the fixed text. Do not delete any information, and do not add any code snippets. Text: {nl} Fixed Text:
@@
      Rewriting perturbed prompts yields only partial recovery: Magicoder‑7B’s pre-processing improves Pass@5 for simple syntactic/noise changes (e.g. formatting fixes) but has almost no effect on deeper semantic edits.
@@
      Overall, automated prompt-fixing yields modest gains limited to trivial noise and often fails on semantic perturbations, making it an unreliable robustness strategy.
@@ 6 Lessons Learned
  • Evaluation benchmarks must move beyond Python. RQ1 reveals: Robust drop varies across languages, whereas prior Python-only work reported uniform patterns. Our work shows that robustness varies substantially across programming languages, whereas prior work Wang et al. (2022) based solely on Python reported relatively consistent robustness patterns across perturbation types within a single language. In contrast, our multilingual results reveal that the magnitude and distribution of robustness drops differ across languages (e.g., Java is more resilient, while C++ degrades more sharply). This comparison indicates that Python-only evaluations may mask language-specific vulnerabilities. These findings motivate the need for community efforts to develop multilingual robustness benchmarks that capture the broader challenges faced by LLMs in real code-generation settings.
@@ 8 Conclusion and Future Work
This paper addresses the demand for a comparative study on the robustness of code generation models across multiple popular programming languages, including Java, C++, and JavaScript. Building upon the foundation of ReCode, which primarily focuses on Python, we extend our work to encompass these languages. Through our research, we adapt 29 distinct types of perturbations from ReCode within four scopes of the prompt. Our investigation yields intriguing findings, which we present in this paper. Notably, we observe that code generation models exhibit varying behaviors across different programming languages and in response to different perturbations. The inherent characteristics of the programming languages have a notable impact on the models’ performance. However, conducting such black box experiments comes with significant computational costs. Therefore, as a future endeavor, we aim to explore white box strategies to enhance the robustness analysis of code generation models.

### v5

@@ Abstract
Large language models have gained significant traction and popularity in recent times, extending their usage to code-generation tasks. While this field has garnered considerable attention, the exploration of testing and evaluating the robustness of code generation models remains an ongoing endeavor. Previous studies have primarily focused on code generation models specifically for the Python language, overlooking other widely-used programming languages. In this work, we conduct a comprehensive comparative analysis to assess the robustness performance of several prominent code generation models and investigate whether robustness can be improved by repairing perturbed docstrings using an LLM. Furthermore, we investigate how their performance varies across different programming languages. To accomplish this, we introduce perturbations in four key areas of the prompt: DocString, function name, syntax, and format. We have compiled and released a dedicated dataset for this purpose. Our results show that all models consistently degrade under perturbations across all three languages, but vary in magnitude depending on the language and perturbation type. Larger model size does not reliably improve robustness, and semantic perturbations prove at least as disruptive as syntactic ones. Our LLM-based docString repair yields only marginal gains for simple perturbations and can degrade performance for semantic ones, highlighting the limits of prompt-level mitigation.
@@ 1 Introduction
Our evaluation of six LLMs across Java, C++, and JavaScript reveals several noteworthy findings. First, robustness drops are consistent across all three languages, but their magnitude varies depending on the language and perturbation type, indicating that robustness is a language-dependent property that cannot be assessed from a single language alone. Second, semantic perturbations prove at least as disruptive as syntactic ones, underscoring the need for diverse perturbation types in robustness benchmarks. Third, model size does not reliably predict robustness. In several cases, larger models are more brittle than smaller ones under semantics-preserving perturbations. Finally, our prompt-repair strategy, using an LLM to reconstruct perturbed docstrings before code generation, yields only marginal gains for simple perturbations and can occasionally degrade performance for semantic ones, suggesting that lightweight prompt-level repairs alone are insufficient to restore robustness.
To guide our investigation, this study addresses the following three research questions:
  1. (1)
    RQ1: How robust are LLMs across different programming languages?
    We evaluate how the performance of six LLMs degrades under semantics-preserving perturbations across Java, C++, and JavaScript, using the EvalPlus-X test suites.
  2. (2)
    RQ2: What are the major factors that contribute to the difference in the robustness of LLMs on different programming languages?
    We investigate which input and output features are most strongly associated with robustness failures, and whether these factors differ across languages.
  3. (3)
    RQ3: Can the robustness drop be mitigated by fixing the perturbed prompts?
    We explore whether using an LLM to automatically repair perturbed docstrings before code generation can recover the lost robustness.
@@ 2.1 Background
In this work, we focus on evaluating the robustness of code generation models. During robustness testing, we apply small, semantics-preserving modifications to the original prompt, for example, changing a single character in a function name, and then ask the model to generate code from this perturbed input. After generation, the correctness of the model’s output is assessed using the same test cases as in the nominal (unperturbed) setting. Following prior work, ReCode (Wang et al., 2022), which studies robustness only in Python, we adapt and extend its 29 perturbation techniques to support C++, Java, and JavaScript.
@@ 2.2 Related Work
Besides, recent studies have explored various aspects of code generation. Such as, Li et al. (2024a) and Li et al. (2024b) have focused on secure code generation, while Ling et al. (2025) investigate social bias in code generation. Lin et al. (2024) examine code generation through the lens of emulating the software development process, and Saha et al. (2024) propose methods for generating code using the source code from other programming languages. More recently, Cheng and Yang (2025) evaluate security aspects in LLM-generated code, Li et al. (2025b, a) investigate automated pipelines for secure code generation, and Rabbi et al. (2025) propose an agentic approach to code translation with specification alignment.
The studies above leave several important questions unaddressed. First, existing robustness evaluations of LLMs for code generation (Wang et al., 2023; Mastropaolo et al., 2023) have focused exclusively on Python, leaving it unclear whether their findings generalize to other widely-used programming languages with distinct syntax and semantics. Second, the benchmarks used in prior robustness work rely on HumanEval, whose test cases have been shown to be insufficient (Liu et al., 2023), potentially masking failures that would be exposed under more rigorous evaluation. Third, while prompt-repair and input-smoothing techniques have been explored in NLP (Ye et al., 2020; Ji et al., 2024), their applicability to code generation robustness remains unexplored. This work directly targets these gaps by extending robustness evaluation to Java, C++, and JavaScript, adopting a stronger test suite based on EvalPlus, and investigating whether LLM-based docstring repair can recover lost robustness.
@@ 3.2 Dataset Preparation
To address the limitations of Wang et al. (2022), which only evaluates code generation robustness in a single programming language (i.e., Python), we extend the evaluation to a multilingual setting by utilizing the HumanEval-X dataset. HumanEval-X, introduced by Zheng et al. (2023), serves as a benchmark for assessing the multilingual capabilities of code generation models. The dataset includes Python, C++, Java, JavaScript, and Go. From the remaining languages, we selected C++, Java, and JavaScript as the most representative and widely used languages in both industry and open-source development, covering diverse language paradigms (compiled vs. interpreted, statically vs. dynamically typed). Go was not selected as it is paradigmatically similar to C++ and Java, also compiled and statically typed, and would therefore contribute limited additional diversity to our evaluation. We also exclude Python from our evaluation since it was already thoroughly studied in prior work, ReCode (Wang et al., 2022). We focused on C++, Java, and JavaScript, which represent diverse language paradigms (compiled vs. interpreted, statically vs. dynamically typed) and are widely used in real-world software development. This selection allows for a balanced and representative evaluation across different programming styles.
@@
The translation was carried out by three researchers, each independently responsible for one target language. We extend this dataset to multilingual by manually rewriting the prompts, canonical solutions, and test cases in C++, Java, and JavaScript through the following steps:
@@
    We find it necessary to update certain canonical solutions because they failed to handle corner cases, inputs at the edges of valid ranges or uncommon conditions that reveal latent bugs. Figures 2 and 8 show two representative failures: in Figure 2 the model produces a wrong answer for a corner input, while in Figure 8 the original solution leads to a timeout. Both examples demonstrate that passing the standard HumanEval-X tests does not guarantee robustness. To mitigate this, we follow the EvalPlus methodology to refine the affected canonical solutions so they correctly handle these edge cases. In total, between 30 and 50 canonical solutions were modified for each of the three languages. Each modified solution was validated by executing it against all converted test cases, and only solutions that passed were retained. After applying those fixes, we generate prompts from the revised solutions for subsequent evaluation.
@@
    We convert the test cases from EvalPlus, originally written in Python, into Java, C++, and JavaScript. During this process, we ensure that data types are transformed into their equivalent counterparts (e.g., converting lists to ArrayList in Java and vector in C++).
@@
    We exclude test cases that rely on Python-specific type flexibility that does not exist in statically typed languages. For example, Python allows heterogeneous lists containing mixed types such as integers and strings, which are not valid in C++ or Java. Similarly, Python supports arbitrarily large integers, whereas Java, C++, and JavaScript are bounded by fixed-size data types. These exclusions are therefore driven by fundamental type system incompatibilities rather than difficulty. The impact of these exclusions is minimal: the original EvalPlus Python benchmark has an average of 764.1 test cases per problem, while our translated versions retain an average of 722.56 (C++), 748.48 (Java), and 709.99 (JavaScript) as stated in  Table3, a reduction of less than 7% in all cases. We therefore do not expect these exclusions to materially affect the difficulty or validity of our benchmark. Throughout this process, we strive to maintain the method signature, including return types and parameter types, identical to those found in the original prompts.
@@ 4.1 Dataset
EvalPlus. EvalPlus (Liu et al., 2023) is an enhanced version of HumanEval, originally developed for Python. It introduces a substantially larger set of corner cases and refined prompts that are better aligned with the associated test cases. In this work, we adapt the EvalPlus test cases for our three target programming languages, following the curation process described in Section 3.2. The prompts are largely consistent with those in HumanEval-X, except in instances where minor modifications were required to accommodate language-specific corner cases.
@@ 4.3 Setup
The same set of hyperparameters is applied across models, chosen through a small grid search on a held-out subset of the evaluation data. The temperature is set to 0.2 to reduce randomness and increase the likelihood of generating syntactically valid and test-passing code. The nucleus sampling parameter (top_p) is set to 0.95 to retain diversity while suppressing low-probability outputs. These values are taken from the ReCode (Wang et al., 2022). The maximum generation length is set to 1536 tokens to prevent truncation. This value is based on our experience that it provides sufficient output while keeping inference time low.
@@ 5.1 RQ1: How robust are LLMs across different programming languages?
Approach: To investigate this research question, we conduct experiments on Java, C++, and JavaScript using our EvalPlus-X derived from EvalPlus (Liu et al., 2023) and HumanEval-X (Zheng et al., 2023). We evaluate six LLMs for code generation: Incoder-1B, Incoder-6B, CodeGen-Multi-2B, CodeGen-Multi-6B, Magicoder-7B, and QwenCode-3.5-6.7B. We evaluate the performance using three metrics: $RP_{s}@k$ measures robustness, $RD_{s}@k$ assesses the decrease in robustness due to changes, and $RR_{s}@k$ captures the relative differences between the original and modified datasets. The perturbations related to the Function Name and the DocString are compared with the nominal prompts (prompts that have only DocString and method signature), while the perturbations related to Syntax and Format are compared with the partial prompts (prompts that have some lines of code along with DocString and method signature).
@@
These behavioral differences across languages can be partly attributed to their type system characteristics. Statistically typed languages such as Java and C++ enforce strict type checking at compile time, meaning that perturbations, particularly syntax-level changes, are more likely to produce compilation errors, directly amplifying the robustness drop. In contrast, JavaScript’s dynamic typing offers more flexibility at runtime, allowing the model to recover from certain surface-level changes without triggering compilation failures. This aligns with our observation that JavaScript tends to show fewer compilation errors and more assertion errors under perturbation, while Java and C++ exhibit the opposite pattern.
@@
    Stronger test cases expose more failures across all languages. From Tables 8 and 9, it is evident that the stronger test cases from EvalPlus reveal substantially more failures in both nominal and perturbed settings, which aligns with prior findings (Liu et al., 2023). The same trend holds for the partial cases, as shown in Tables 10 and 11. However, the robustness drop remains roughly consistent regardless of the absolute performance levels.
@@ 5.3 Can the robustness drop be mitigated by fixing the perturbed prompts?
Approach: We focus on the DocString perturbation dataset for this experiment. Our approach is inspired by prior work in natural language processing, where LLMs are used to correct typographical and grammatical errors in English sentences (Coyne et al., 2023). Specifically, we select all ten DocString-perturbed subsets from our dataset and employ the Magicoder-7B model to fix the prompts. We choose Magicoder because it is the only instruction-tuned model among our six evaluated models capable of reliably following natural language repair instructions. The remaining five models: Incoder-1B, Incoder-6B, CodeGen-2B-Multi, CodeGen-6B-Multi, and QwenCode, are primarily optimized for code generation and tend to produce code snippets rather than corrected natural language text when given repair prompts, making them unsuitable for this task.
@@
Fix the grammatical errors, spelling, and typos in the following text and reply with the fixed text. Do not delete any information, and do not add any code snippets. Text: {nl} Fixed Text:
@@
The limited recovery for semantic perturbations can be explained by the nature of the repair prompt, which instructs the model to fix grammar, spelling, and typos. Perturbations such as BackTranslation and Synonym-Substitution produce docstrings that are already grammatically correct and fluent, leaving the repair model with no clear signal to identify or correct. In some cases, the model rewrites the docString in a way that diverges further from the original intent, explaining the occasional performance degradation visible in Figure 10. Future approaches could explore context-aware repair strategies that use the function signature as an anchor to guide docString reconstruction, or employ retrieval-augmented generation to better align the repaired docString with the intended functionality. For other perturbation types, function name perturbations could, in principle, benefit from a similar word-level repair strategy, while syntax and format perturbations target the code structure directly and would require fundamentally different, code-specific repair strategies.
@@
    Prompt repair yields only partial recovery, and its effectiveness is tied to perturbation type: it helps for surface-level changes such as whitespace or character modifications, but has little effect on deeper semantic edits where the perturbed text is already natural and fluent.
@@
    The repair strategy is ineffective for semantic perturbations because they produce docstrings that are already grammatically valid, providing no signal for the repair model to act on. This is the key reason why BackTranslation and SynonymSubstitution show little or no recovery, and sometimes degrade performance. Future work should explore context-aware repair strategies, for instance, using the function signature as an anchor to reconstruct a semantically aligned docString.
Extending the repair strategy to other perturbation types remains an open challenge. Syntax and format perturbations target code structure directly and would require code-specific repair strategies such as AST-based transformations. For function name perturbations, our natural language sentence repair approach does not directly apply, as function names consist of one or more words without sentence structure. Repairing them would require parsing the name into constituent words and replacing corrupted ones with the closest valid alternatives, a fundamentally different strategy that, given the limited context of a function name, remains an open direction for future work.
@@ 6 Lessons Learned
  • Evaluation benchmarks must move beyond Python. RQ1 reveals: Robust drop varies across languages, whereas prior Python-only work reported uniform patterns. Our work shows that robustness varies substantially across programming languages, whereas prior work (Wang et al., 2022) based solely on Python reported relatively consistent robustness patterns across perturbation types within a single language. In contrast, our multilingual results reveal that the magnitude and distribution of robustness drops differ across languages (e.g., Java is more resilient, while C++ degrades more sharply). This comparison indicates that Python-only evaluations may mask language-specific vulnerabilities. These findings motivate the need for community efforts to develop multilingual robustness benchmarks that capture the broader challenges faced by LLMs in real code-generation settings.
@@ 7 Threats to Validity
In addition, the benchmarks used in this study, HumanEval-X and EvalPlus, primarily consist of small, self-contained algorithmic tasks. While these benchmarks are widely adopted in the code generation community and enable controlled, reproducible evaluation, they may not fully represent the complexity of real-world software development tasks, which often involve larger codebases, complex dependencies, and domain-specific requirements. As a result, the robustness findings reported in this study may not directly generalize to more complex real-world scenarios. Evaluating robustness on more representative benchmarks covering diverse and complex programming tasks remains an important direction for future work.
Another external threat is that the six models evaluated in this study are all open-source and locally deployable. Widely used closed-source models such as GPT-4, CodeLlama, Claude and Gemini are not included in this study. The excluded models differ in architecture, training data, and scale, and may exhibit different robustness patterns under the same perturbations. In particular, larger instruction-tuned or RLHF-trained models may be more resilient to surface-level perturbations due to their stronger natural language understanding. As a result, the conclusions drawn from our six models may not fully generalize to closed-source or larger-scale models, and we acknowledge this as a limitation of our study.
In this study each prompt is perturbed by a single perturbation at a time, following the evaluation design of ReCode (Wang et al., 2022). In practice, a developer may inadvertently introduce multiple surface-level changes to a prompt simultaneously, for example, a typo in the function name combined with a paraphrased docString. Applying multiple perturbations simultaneously could compound the robustness drop beyond what is observed in single-perturbation settings, and the interaction effects between perturbation types remain an open question for future work.
Construct Validity: The perturbations used in this study target DocString, function name, syntax, and format, all of which are surface-level or natural language modifications (e.g., variable names). Our study does not include type-related perturbations, such as changes to type annotations, variable types, or type casts, which may affect statically typed languages like Java and C++ differently from dynamically typed ones like JavaScript. The absence of such perturbations may limit the completeness of our robustness evaluation with respect to language-specific type system characteristics. We consider the design and evaluation of type-aware perturbations as a direction for future work.
@@ 8 Conclusion and Future Work
This paper addresses the demand for a comparative study on the robustness of code generation models across multiple popular programming languages, including Java, C++, and JavaScript. Building upon the foundation of ReCode (Wang et al., 2022), which primarily focuses on Python, we extend the robustness evaluation to a multilingual setting by constructing EvalPlus-X, a new benchmark derived from HumanEval-X and EvalPlus with substantially more rigorous test cases. We evaluate six LLMs, Incoder-1B, Incoder-6B, CodeGen-2B-Multi, CodeGen-6B-Multi, Magicoder-7B, and QwenCode-2.5-7B, under 29 distinct semantic-preserving perturbations across four prompt scopes: DocString, function name, syntax, and format.
Our investigation yields several key findings. All models consistently degrade under perturbations across all three languages, but the magnitude of robustness drops varies depending on the language and perturbation type, confirming that robustness is a language-dependent property that cannot be adequately assessed from a single language alone. Semantic and format perturbations prove at least as disruptive as syntactic ones, underscoring the need for diverse perturbation types in robustness benchmarks. Larger model size does not reliably predict robustness; in several cases, larger models are more brittle than smaller ones under semantics-preserving perturbations. Our feature-level analysis (RQ2) further reveals that the factors driving robustness drops differ across languages, with DocString and function name perturbations most strongly associated with failures in Java and JavaScript, while C++ is more sensitive to syntax-level perturbations. Our LLM-based docstring repair strategy (RQ3) yields only marginal gains for surface-level perturbations and is ineffective for semantic ones, highlighting the limits of lightweight prompt-level mitigation.
Several directions remain open for future work. This study evaluates each perturbation type in isolation, following ReCode (Wang et al., 2022), which allows the individual effect of each perturbation to be isolated. However, real-world scenarios may involve multiple simultaneous changes to a single prompt, and prior NLP robustness work has shown that compound perturbations can reveal interaction effects beyond what single perturbations expose (Dong et al., 2023; Gupta et al., 2024). Extending our framework to support combined perturbations across multiple scopes represents a promising direction. Beyond that, evaluating robustness using white-box strategies, which provide access to model internals, could offer deeper insights beyond the black-box setting adopted here. Extending the evaluation to broader and more complex benchmarks, additional programming languages, and larger closed-source models would further strengthen the generalizability of the findings.
Acknowledgement
This research was supported by the Fonds de recherche du Québec (Grant No.2024-NOVA346499)(Fonds de recherche du Québec, 2024), Natural Sciences and Engineering Research Council of Canada (NSERC) through the Alliance, Grant (Grant No.586838-23), the NSERC Discovery Grant (Grant No. RGPIN-2019-07007 and Grant No. DGECR-2019-00464), and NSERC CREATE Grant (Grant No.555406-2021). We gratefully acknowledge the support of all funding agencies.